40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri May 13 14:26:00 2022, JC, Update, Alignment, LO2 Beam 
    Reply  Fri May 13 18:42:13 2022, Paco, Update, Alignment, ITMX and ITMY sat amp failures 
Message ID: 16851     Entry time: Fri May 13 14:26:00 2022     Reply to this: 16852
Author: JC 
Type: Update 
Category: Alignment 
Subject: LO2 Beam 

[Yehonathan, JC]

Yehonathan and I attempted to align the LO2 beam today through the BS chamber and ITMX Chamber. We found the LO2 beam was blocked by the POKM1 Mirror. During this attempt, I tapped TT2 with the Laser Card. This caused the mirror to shake and dampen into a new postion. Afterwards, when putting the door back on ITMX, one of the older cables were pulled and the insulation was torn. This caused some major issues and we have been able to regain either of the arms to their original standings.

ELOG V3.1.3-